OpenStudy (anonymous):

Precalculus Piecewise Functions Help? Determine which is the graph of the function of f(x)= {3x+2 if x<5} {-3x-2 if x>=5}

OpenStudy (anonymous):

@pooja195 could you help me solve this? I'm not sure where to start.

OpenStudy (anonymous):

draw a line y=3x+2 for x<5 and another line y=-3x-2 for x>5

OpenStudy (anonymous):

Okay one minute let me draw it. Thank you so much!

OpenStudy (anonymous):

Okay, but also how do I draw those lines on a graph? Isn't there a certain method that I need to do in order to graph the lines? @surjithayer

OpenStudy (anonymous):

when x=0,y=3*0+2=2 when x=4,y=3*4+2=14 draw a st. line through (0,2) and (4,14)

OpenStudy (anonymous):

Okay, but where are you getting these numbers from my equation? Sorry I'm just a bit confused.

OpenStudy (anonymous):

These are my graphs that I have to choose from.

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you. So this is how my graph should look?

OpenStudy (anonymous):

now you can draw y=-3x-2 for x>5 when x=6 ,y=-3*6-2=-20 x=8,y=-3*8-2=-26 now draw aline through (6,-20) and (8,-26)

OpenStudy (anonymous):

here x>=5

OpenStudy (anonymous):

Okay so I'm drawing another line on the graph with the given points correct?

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay thank you